RHRS magnets is turned off the this shift. Detector checks only on left arm.
22:00 Compton tuning done for tonight. Continueing to detector check on HRSs
22:10 Request beam position to be tuned to be close to the one on white borad : BPMA = (-.35, 1.8); BPMB = (-1.3, 2.1)
22:39 beam position is close to white board one. BPMA = (-.42, 1.7); BPMB = (-1.3, 2.5). Requesting a harp scan.
22:50 Harp result posted. width ~100um. BeO Check OK.
23:10 1uA beam on Carbon Optics. Run 25181